# Augmenting Neural Networks-based Model Approximators in Robotic Force-tracking Tasks

### v1

@@ 2.2 Direct Force Controller
where, if $m=6$, $\bm{\Gamma}=\operatorname{diag}(\gamma_{x},\gamma_{y},\gamma_{z},\gamma_{\phi},\gamma_{\theta},\gamma_{\psi})$ is the task specification matrix [Khatib, 1987], with $\gamma_{i}=1$ if the $i$-th direction is subject to force control, 0 otherwise. $\bm{x}_{f}\in\mathbb{R}^{m}$ is the force controller output, while $\bm{x}_{r}\in\mathbb{R}^{m}$ is the reference pose, tracked when $\gamma=0$. $\bm{K}_{P},\bm{K}_{I}\in\mathbb{R}^{m\times m}$ are the proportional and integral gains of the controller, and $\bm{\Delta}\bm{h}=\bm{h}_{r}-\bm{h}_{e}\in\mathbb{R}^{m}$ is the error between the reference wrench to be exerted $\bm{h}_{r}\in\mathbb{R}^{m}$ and the actual exerted wrench $\bm{h}_{e}\in\mathbb{R}^{m}$.
@@ 3.2 Experiment I: Model Approximator Validation
For a fair comparison with [Petrone et al., 2025], we base the FFNNs structure on the findings therein. Linear type layers are used, while the activation function is of the ReLU type [Agarap, 2019]. The optimized configuration of the base estimator that ensures enhanced inference time without compromising prediction accuracy uses $N=3$ FFNNs, configured with the parameters of Table 2. Adam [Kingma and Ba, 2015] refers to Adaptive Moment Estimation, an optimizer used for NN regression tasks. All FFNNs used have the same configuration.
@@ 3.2.3 Model Approximators Comparison
This qualitative analysis is quantitatively confirmed by Table 3, which also reports the numerical improvement factor $\eta$ of DMA over SMA. As $v$ increases, the latter outperforms the former by up to 454%, in terms of average RMSE recorded across trajectories at the same velocity.
@@ 3.3 Experiment II: Control Algorithm Validation
As can be seen from Fig. 5, the average force-tracking RMSE for VAICAM is approximately constant for $v\in\left[0,0.35\right]\,$\mathrm{m}\text{/}\mathrm{s}$\$, and starts increasing for $v>$0.35\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$. Thanks to the exploitation of the novel DMA, VAICAM outperforms both ORACLE and the base DFC. A sample trajectory is considered in Fig. 6, visually showing that ORACLE’s uncertainty at moderate velocities (due to the limited SMA) translates in worse force-tracking capabilities. Furthermore, Fig. 5 shows that VAICAM also enhances the DFC performance, although with a lower improvement factor compared to ORACLE. Lastly, it is worth noticing that VAICAM’s RMSE standard deviation is lower than DFC’s, thus manifesting a superior robustness w.r.t. the experimental conditions.
@@
To confirm the improvement of VAICAM over both DFC and ORACLE, Table 4 displays the numerical results by reporting the average RMSE of these 3 strategies for each velocity. As the improvement factor $\eta$ confirms, VAICAM surpasses ORACLE by up to 127% at $v=$0.5\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$, and outperforms DFC by up 31% at $v=$0.35\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$.

### v3

@@ 2.2 Direct Force Controller
where, if $m=6$, $\bm{\Gamma}=\operatorname{diag}(\gamma_{x},\gamma_{y},\gamma_{z},\gamma_{\phi},\gamma_{\theta},\gamma_{\psi})$ is the task specification matrix [Khatib, 1987], with $\gamma_{i}=1$ if the $i$-th direction is subject to force control, 0 otherwise. $\bm{x}_{f}\in\mathbb{R}^{m}$ is the force controller output, while $\bm{x}_{r}\in\mathbb{R}^{m}$ is the reference pose, whose $i$-th component is tracked when $\gamma_{i}=0$. $\bm{K}_{P},\bm{K}_{I}\in\mathbb{R}^{m\times m}$ are the proportional and integral gains of the controller, and $\bm{\Delta}\bm{h}=\bm{h}_{r}-\bm{h}_{e}\in\mathbb{R}^{m}$ is the error between the reference wrench to be exerted $\bm{h}_{r}\in\mathbb{R}^{m}$ and the actual exerted wrench $\bm{h}_{e}\in\mathbb{R}^{m}$.
@@ 3.2 Experiment I: Model Approximator Validation
For a fair comparison with [Petrone et al., 2025], we base the FFNNs structure on the findings therein. Specifically, all FFNNs used share the same configuration in terms of depth (number of hidden layers), width (number of neurons per layer) and learning algorithm parameters, as listed in Table 2. Linear type layers are used, while the activation function is of the ReLU type [Agarap, 2019]. The optimized configuration of the base estimator that ensures enhanced inference time without compromising prediction accuracy uses $N=3$ FFNNs. Adam [Kingma and Ba, 2015] refers to Adaptive Moment Estimation, an optimizer used for NN regression tasks, while “Fusion” indicates how a single output is retrieved from the ensemble, i.e. the arithmetic average across the independent network estimations is computed.
@@ 3.2.3 Model Approximators Comparison
This qualitative analysis is quantitatively confirmed by Table 3, which also reports the numerical improvement factor $\eta$ of DMA over SMA. As $v$ increases, the former outperforms the latter by up to 454%, in terms of average RMSE recorded across trajectories at the same velocity.
@@ 3.3 Experiment II: Control Algorithm Validation
As can be seen from Fig. 5, the average force-tracking RMSE for VAICAM is approximately constant for $v\in\left[0,0.35\right]\,$\mathrm{m}\text{/}\mathrm{s}$\$, and starts increasing for $v>$0.35\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$. Thanks to the exploitation of the novel DMA, VAICAM outperforms both ORACLE and the base DFC. A sample trajectory is considered in Fig. 6, visually showing that ORACLE’s uncertainty at moderate velocities (due to the limited SMA) translates into worse force-tracking capabilities. Furthermore, Fig. 5 shows that VAICAM also enhances the DFC performance, although with a lower improvement factor compared to ORACLE. Lastly, it is worth noticing that VAICAM’s RMSE standard deviation is lower than DFC’s, thus manifesting a superior robustness w.r.t. the experimental conditions.
@@
To confirm the improvement of VAICAM over both DFC and ORACLE, Table 4 displays the numerical results by reporting the average RMSE of these 3 strategies for each velocity. As $\eta$ confirms, VAICAM surpasses the DFC by up to 31% (at $v=$0.35\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$), while the improvement over ORACLE at maximum experimental speed ($v=$0.5\text{\,}\mathrm{m}\text{/}\mathrm{s}$\$) reaches 127%.